Cryptology Assignment 1

Due Monday, March 4, 1996. Do problems 1.8, 1.11, 2.1, and 2.4 in the textbook.

Problems

To be discussed Wednesday, February 28, 1996. Do problems 2.3, 2.7 and 2.8 in the textbook.